Please help me solve with quadratic formula. 6x^2=-12x-1. Thank you. Also website/ computer malfunction.. i cannot see topics on the dragdown so the topic will be random.
Answer by stanbon(57385)

Please help me solve with quadratic formula. 6x^2=-12x-1. Thank you. Also website/ computer malfunction.. i cannot see topics on the dragdown so the topic will be random.
----------------------------------
6x^2=-12x-1
---
6x^2+12x+1 = 0
---
x = [-12 +- sqrt(144-4*6*1)]/12
----
x = [-12 +- sqrt(4*30)]/12
---
x = [-1 +-(1/6)sqrt(30)]
===============================
Cheers,
Stan H.
=============